OpenStudy (anonymous):

athletes from both Northville High School and southfield High school rode buses to get to a regional track meet westville. The northville team traveled x miles, and the southfield team traveled (3x - 12) miles. The total traveled by the two teams wwere 64 miles. How far did the Soouthfield team travel to get to the track meet?

OpenStudy (anonymous):

x + 3x - 12 = 64, find x Southfield team traveled (3x-12) miles

OpenStudy (anonymous):

or 27!

OpenStudy (lolo):

what?

OpenStudy (yuki):

x and 3x-12 are the total miles that they both traveled, since it is 64 miles x + (3x-12) = 64 if you solve for it, 4x = 76 thus x = 19

OpenStudy (anonymous):

whoops, subtracted instead of added...

OpenStudy (yuki):

since the other traveled 3x-12, 3*19-12 = 57 - 12 = 35

OpenStudy (anonymous):

I GOT 25.3 LOL

OpenStudy (yuki):

oops, 45

OpenStudy (yuki):

cause 45 + 19 = 64 that makes more sense

OpenStudy (lolo):

which 1 Is the right answer if ANSWER THE QUESTION !!!!!!!